A tank holds 80 liters of a chemical solution. Currently the solution has a strength of 30%. How much of this solution must be drained and replaced with a 70% solution to get a strength of 40%?
Found 2 solutions by stanbon, scott8148:
Answer by stanbon(48558)

A tank holds 80 liters of a chemical solution. Currently the solution has a strength of 30%. How much of this solution must be drained and replaced with a 70% solution to get a strength of 40%?
-------------
30% solution DATA:
Amt = 80 L ; amt of active ingredient = 0.3*80 = 24 L
----------
70% solution DATA:
Amt = x L ; amt of active ingredient = 0.7x L
-------------
40% remainder DATA:
Amt = (80) L ; Amt of active ingredient = 0.4(80) = 32 L
-------------------
EQUATION:
active - active = active
24 + 0.7x = 32
0.7x = 8
x= 11.43 L
-------------------
Cheers,
Stan H.

Answer by scott8148(5891)

When the tank is filled with 40% solution, there will be 32 liters of chemical in the solution (40% of 80 liters or .4 X 80). y liters of 30% solution must be drained and replaced with 70% solution.

(chemical left in tank) + (chemical added) = 32 liters

{(80-y) X .30} + (y X .70) = 32

(24-.3y) + .7y = 32

.4y = 8

y = 20 liters